OpenStudy (anonymous):

You are selling tickets for a high school basketball game. Student tickets cost $3 and general admission tickets cost $5. You sell 350 tickets and collect $1,450. How many student tickets did you sell?

OpenStudy (anonymous):

This is a system of equations with 2 variables. What you first need to figure out is what is x and what is y. What are the two things that you don't know?

OpenStudy (anonymous):

need 2 know student ticket #

OpenStudy (anonymous):

What is the other item that you don't know? There are two. You guessed one, number of student tickets sold. What is your other unknown?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

3x + 5y = 1450 -3x -3x 5y = (1450 - 3x) idk I'm pretty sure thats the first step

OpenStudy (anonymous):

There are two types of tickets being sold: Student tickets and general admission tickets.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Now, what are the two equations that you will have that describe this situation. One equation will be the total number of tickets sold, which is a combination of student tickets and general admissions tickets. What is the other equation?

OpenStudy (anonymous):

sorry had 2 g 2 restroom umm is it 3*x+5*x=1,450 @dcabeche

OpenStudy (anonymous):

Yes, your two equations should be \[x + y = 350\] and \[3x +5y = 1450\]

OpenStudy (anonymous):

oh okay thks a lot that really helped :)

OpenStudy (anonymous):

200 standard 5 dollar tickets, 150 three dollar tickets. Lots of trial and error because I didn't want to make a system with mental math.